# A halo model approach for mock catalogs of time-variable strong gravitational lenses

### v4

@@ Acknowledgements.
Author contributions: KTA mainly developed the SL-Hammocks code, conducted most of the analysis in the paper, and wrote the main body of the paper. MO conceived of and initiated the project and contributed to the development of the SL-Hammocks code, designing the validation tests, interpretations of results, and paper writing. SB provided a code to compute velocity dispersions of galaxies for validations, and contributed to the development and validations of the SL-Hammocks code. NK and PJM also contributed to the development and validations of the SL-Hammocks code. CL and AM provided suggestions on many aspects of the analysis and the write-ups. We thank anonymous referees for constructive comments. This work was supported by JSPS KAKENHI Grant Numbers JP23K22531, JP20H05856, JP22K21349. The work of PJM was supported by the U.S. Department of Energy under contract number DE-AC02-76SF00515. This project has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodovska-Curie grant agreement No 101105725. This paper has undergone internal review in the LSST Dark Energy Science Collaboration. We would like to thank internal reviewers Cameron Lemon and Anupreeta More for their role. The DESC acknowledges ongoing support from the Institut National de Physique Nucléaire et de Physique des Particules in France; the Science & Technology Facilities Council in the United Kingdom; and the Department of Energy, the National Science Foundation, and the LSST Corporation in the United States. DESC uses resources of the IN2P3 Computing Center (CC-IN2P3–Lyon/Villeurbanne - France) funded by the Centre National de la Recherche Scientifique; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231; STFC DiRAC HPC Facilities, funded by UK BEIS National E-infrastructure capital grants; and the UK particle physics grid, supported by the GridPP Collaboration. This work was performed in part under DOE Contract DE-AC02-76SF00515. This work used the following packages; NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), COLOSSUS (Diemer, 2018), glafic (Oguri, 2010, 2021), joblib (Joblib Development Team, 2020), Astropy (Astropy Collaboration et al., 2022).

### v5

@@ Acknowledgements.
Author contributions: KTA mainly developed the SL-Hammocks code, conducted most of the analysis in the paper, and wrote the main body of the paper. MO conceived of and initiated the project and contributed to the development of the SL-Hammocks code, designing the validation tests, interpretations of results, and paper writing. SB provided a code to compute velocity dispersions of galaxies for validations, and contributed to the development and validations of the SL-Hammocks code. NK and PJM also contributed to the development and validations of the SL-Hammocks code. CL and AM provided suggestions on many aspects of the analysis and the write-ups. We thank Kai Liao for pointing out a bug in calculating lensing by satellite galaxies, which is now fixed. We thank anonymous referees for constructive comments. This work was supported by JSPS KAKENHI Grant Numbers JP23K22531, JP20H05856, JP22K21349. The work of PJM was supported by the U.S. Department of Energy under contract number DE-AC02-76SF00515. This project has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodovska-Curie grant agreement No 101105725. This paper has undergone internal review in the LSST Dark Energy Science Collaboration. We would like to thank internal reviewers Cameron Lemon and Anupreeta More for their role. The DESC acknowledges ongoing support from the Institut National de Physique Nucléaire et de Physique des Particules in France; the Science & Technology Facilities Council in the United Kingdom; and the Department of Energy, the National Science Foundation, and the LSST Corporation in the United States. DESC uses resources of the IN2P3 Computing Center ($C$-IN2P3–Lyon/Villeurbanne - France) funded by the Centre National de la Recherche Scientifique; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231; STFC DiRAC HPC Facilities, funded by UK BEIS National E-infrastructure capital grants; and the UK particle physics grid, supported by the GridPP Collaboration. This work was performed in part under DOE Contract DE-AC02-76SF00515. This work used the following packages; NumPy (2020Natur.585..357H), SciPy (2020SciPy-NMeth), Matplotlib (2007CSE.....9...90H), COLOSSUS (2018ApJS..239...35D), glafic (2010PASJ...62.1017O; 2021PASP..133g4504O), joblib (joblib), Astropy (2022ApJ...935..167A).